OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -1/4 and y-intercept (0, -2) and graph the line.

OpenStudy (lgbasallote):

hint: the slope intercept form is y= mx + b where m is slope and b is the y-intercept so if you have a slope of 2 and a y-intercept of (0,1) then the equation will be y = 2x + 1 does that help?

OpenStudy (anonymous):

I'm so confused.

OpenStudy (lgbasallote):

with what?

OpenStudy (anonymous):

Everything, I don't understand the slope and y intercept at all.

OpenStudy (lgbasallote):

slope and y-intercept form is y = mx + b m is slope b is y-intercept do you get that?

OpenStudy (anonymous):

yes that I do get

OpenStudy (lgbasallote):

so in your problem what is the slope

OpenStudy (anonymous):

m=-1/4?

OpenStudy (lgbasallote):

yes. and what is the y-intercept?

OpenStudy (anonymous):

y=-2

OpenStudy (lgbasallote):

dont you mean b = -2 b is the letter for y-intercept

OpenStudy (lgbasallote):

so substitute m=-1/4 and b=-2 into y = mx + b

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

ok so....it's y = mx + b? Is that my answer?

OpenStudy (lgbasallote):

no. SUBSTITUTE m= -1/4 and b = -2

OpenStudy (lgbasallote):

look at the example i wrote in my first comment. that's how you should go

OpenStudy (anonymous):

So the slope of 2 is that refering to the -2 in parentheses? I'm sorry for all the questions.

OpenStudy (lgbasallote):

nahh just keep the questions coming lol

OpenStudy (anonymous):

y = 2x + 1

OpenStudy (lgbasallote):

actually that's the answer for my EXAMPLE...not this problem

OpenStudy (anonymous):

oh ok lol

OpenStudy (lgbasallote):

notice how the given are different...i said the slope is 2 and the y-intercept is (0,1) in your case the slope is -1/4 and y-intercept is (0,-2)

OpenStudy (anonymous):

y = -1/4 + -2

OpenStudy (anonymous):

-1/4x

OpenStudy (lgbasallote):

so...which one is your answer?

OpenStudy (anonymous):

y = -1/4x + -2

OpenStudy (lgbasallote):

right.. or just write it as y = -1/4 x - 2

OpenStudy (lgbasallote):

that's the final answer now

OpenStudy (anonymous):

(Whew) got it....thanks so much! I wish I could give you 2 medals for putting up with me lol

OpenStudy (lgbasallote):

haha nahh it's fine. i like it when people ask about the process :)) anyway i have to go now so good luck

OpenStudy (anonymous):

Thanks, good night